Derek plans to retire on his 65th birthday. However, he plans to work part-time until he turns 72.00. During these years of part-time work, he will neither make deposits to nor take withdrawals from his retirement account. Exactly one year after the day he turns 72.0 when he fully retires, he will wants to have $3,492,681.00 in his retirement account. He he will make contributions to his retirement account from his 26th birthday to his 65th birthday. To reach his goal, what must the contributions be? Assume a 10.00% interest rate.
Answer #1

Rate of interest = 10%

n = 72 - 65 = 7 years

Value of money desired at the age of 65, at the time he is 65 = $3,492,681/(1+r)n

= $3,492,681/(1+0.1)7

= $3,492,681/(1.1)7

= $3,492,681/1.949

= $1,792,037.46

Yearly regular deposits = R

Number of years of regular deposits(m) = 65 -26 = 39 years

Future value of investment(S) = R[{(1+r)m-1}/r] (1+r)

1792037.46 = R[{(1.1)39- 1}/0.1](1.1)

1792037.46 = R[{41.15-1}/0.1](1.1)

1792037.46 = R(40.15*11)

1792037.46 = R(441.65)

R = $4057.60 Answer

To reach his goal, $4057.60 contributions must be made by him yearly.